# Optimal Inapproximability of Promise Equations

### v2

@@ 1 Introduction
In this work, we strengthen the optimal inapproximability results for $\operatorname{3-LIN}$ over finite groups by establishing NP-hardness of beating the random assignment threshold even if the instance is almost satisfiable in an arbitrarily more restrictive setting. Formally, this is captured by fixing (not one but) two groups and a homomorphism between them, following the framework of promise CSPs [AGH17, BBKO21]. In detail, (decision) promise CSPs [BBKO21] can be seen as a qualitative form of approximation: Each constraint comes in two forms, a strong one and a weak one. The promise is that there is a solution satisfying all constraints in the strong form while the (potentially easier) goal is to find a solution satisfying all constraints in the weak form. An example of a strong vs. weak constraint on the same, say Boolean, domain is $1$-in-$3$ vs NAE, where the former is $\{(0,0,1),(0,1,0),(1,0,0)\}$ and the latter is $\{(0,0,1),(0,1,0),(1,0,0),(1,1,0),(1,0,1),(0,1,1)\}$. NAE is weaker as the relation contains more tuples. While these two constraint relations capture the well-known NP-hard problems of 1-in-3-SAT and Not-All-Equal-SAT respectively [Sch78], finding an NAE-assignment turns out to be doable in polynomial time under the promise that a 1-in-3-assignment exists [BG21]! [^1] The algorithm involves solving an instance of $\operatorname{3-LIN}$ over the integers and rounding positive integers to $1$ and non-positive integers to 00, demonstrating the importance of $\operatorname{3-LIN}$ among promise CSPs. For constraints on different domains, the notion of strong vs. weak constraint is captured by a homomorphism between the (sets of all) constraint relations; in the example above, the homomorphism is just the identity function. The exact solvability of $\operatorname{3-LIN}$ in the promise setting was resolved in [LŽ25].
@@
Recent work of Barto et al. [BBK+24] initiated an algebraic investigation of (quantitative) approximation of promise CSPs. In the context of $\operatorname{3-LIN}$, here are two simple examples captured by this framework. First, let $\mathscr{G}$ be a group and $\mathscr{H}$ be a subgroup of $\mathscr{G}$. Given an almost-satisfiable system over the subgroup $\mathscr{H}$, maximise the number of satisfied equations over $\mathscr{G}$. Our results imply that beating the random assignment over $\mathscr{H}$ is NP-hard. In the second example, consider a group $\mathscr{G}$, a normal subgroup $\mathscr{H}$, and an almost-satisfiable system over $\mathscr{G}$. The goal this time is to maximise the number of satisfied equations in the system over the quotient $\mathscr{G}/\mathscr{H}$. Our results show that doing better than the random assignment over $\mathscr{G}/\mathscr{H}$ is NP-hard. More generally, going beyond subgroups and quotients of a given group, we fix two groups $\mathscr{G}_{1}$ and $\mathscr{G}_{2}$ and a group homomorphism $\varphi$ from a subgroup $\mathscr{H}_{1}$ of $\mathscr{G}_{1}$ to a subgroup $\mathscr{H}_{2}$ of $\mathscr{G}_{2}$ with the property that $\varphi$ extends to a group homomorphism from $\mathscr{G}_{1}$ to $\mathscr{G}_{2}$. Given an almost-satisfiable system of equations over $\mathscr{G}_{1}$ with constants in $\mathscr{H}_{1}$, the goal is to maximise the number of satisfied equations over $\mathscr{G}_{2}$ where the constants are interpreted in $\mathscr{H}_{2}$ via $\varphi$. Our main result establishes that doing better than the random assignment over $\mathscr{H}_{2}$ is NP-hard, cf. Theorem 3.1. Thus we give an optimal inapproximability result for a natural and fundamental fragment of promise CSPs, systems of linear equations.
@@
The structure of the paper is as follows. In Section 2 we set the notation and present the necessary technical background on Fourier analysis over non-Abelian groups. Section 3 presents the main results with discussion and also provides a sketch of the main proof: In Section 3.1 we present the reduction from Gap Label Cover to Promise $\operatorname{3-LIN}$, and in Section 3.2 we give an overview of the techniques used in the analysis of this reduction and of the main challenges that arise in extending previous work to the promise setting. Section 4 then gives all technical details of the proof of the main result, with the completeness analysis in Section 4.1 and the soundness analysis in Section 4.2. Finally, in Section 5 we relate our results to a recent theory of Barto et al. [BBK+24], who developed a systematic approach to study (in)approximability of promise CSPs, which includes approximability of promise linear equations, from the viewpoint of universal algebra. In particular, we show that the proof of our main result (Theorem 3.1) implies that the collection of symmetries222called the valued minion of plurimorphisms in [BBK+24]. of Promise $\operatorname{3-LIN}$ can be mapped homomorphically to the collection of symmetries of Gap Label Cover, a condition that, based on the algebraic theory from [BBK+24], is known to guarantee NP-hardness of the former problem.
@@ 2 Background
[…]
@@ 2.1 Fourier Analysis over non-Abelian Groups
[…]
@@
Let $N$ be a finite set. Given a pair of functions function $F,H:\mathscr{G}\to\mathbb{C}^{N\times N}$, we define their convolution $F*H$ by
@@ 2.1.1 Fourier Analysis over Direct Products
Let $\mathscr{G}$ be a finite group, $D$ a finite set, and $10<\epsilon<$ 1 a real number. Let $\bm{\nu}=(\nu_{d})_{d\in D}$ be a random element from $\mathscr{G}^{D}$ chosen as follows. Independently for each $d\in D$, the element $\nu_{d}$ equals $1_{\mathscr{G}}$ with probability $1-\epsilon$, and is picked uniformly at random from $\mathscr{G}$ otherwise. Let $F:\mathscr{G}^{D}\to\mathbb{C}$ and define $H(\mathbf{a}):=\mathbb{E}_{\bm{\nu}}[F(\mathbf{a}\cdot\bm{\nu})]$. Then, for each $\rho\in\widehat{\mathscr{G}^{D}}$ and $i,j\in N_{\rho}$, we have
@@ 3 Overview of Results
Let $(\mathscr{G}_{1},\mathscr{G}_{2},\varphi)$ be a cubic template and $10<s\leq c<$ 1. Then, $\operatorname{3-LIN}(\mathscr{G}_{1},\mathscr{G}_{2},\varphi,c,s)$ is tractable if $s\leq 1/|\textnormal{Im}(\varphi)|$ and NP-hard otherwise.
@@
Given a weighted system of equations over $(\mathscr{G}_{1},\mathscr{G}_{2},\varphi)$, consider its set of unsatisfiable equations. Since $\varphi$ extends to a full homomorphism, if the total weight of the set of unsatisfiable equations is more than $1-c$, then the instance cannot be $c$-satisfiable in $\mathscr{G}_{1}$, hence, reject. Otherwise, the random assignment over $\textnormal{Im}(\varphi)$ satisfies at least a $1/|\textnormal{Im}(\varphi)|$-fraction of the satisfiable equations over $\mathscr{G}_{2}$, which is at least a $c/|\textnormal{Im}(\varphi)|$-fraction of the entire system. It is a simple corollary of Theorem 3.1 that this algorithm is optimal for non-cubic groups, leading to the following result. Details are deferred to Appendix A.
@@
Let $(\mathscr{G}_{1},\mathscr{G}_{2},\varphi)$ be a non-cubic template and $10<s\leq c<$ 1. Then, $\operatorname{3-LIN}(\mathscr{G}_{1},$ $\mathscr{G}_{2},\varphi,c,s)$ is tractable if $s/c\leq 1/|\textnormal{Im}(\varphi)|$ and NP-hard otherwise.
@@ 3.2 Proof Outline
$\left|\mathrm{tr}\,\mathbb{E}\left[\left(\sum_{\tau\in\widehat{\mathscr{G}_{1}^{E}},\tau\neq 1}\sum_{s,t\in N_{\tau}}\dim_{\tau}\widehat{\mathcal{A}}(\tau_{s,t})\tau_{s,t}({\mathbf{a}})\right)\times\right.\right.$
@@ 4.1 Completeness
Let $f_{U}:U\to D$, $f_{V}:V\to E$ be assignments satisfying the original GLC instance $\Sigma$ that is to say, these assignments are such that $\pi_{uv}(f_{U}(u))=f_{V}(v)$ for each edge $\pi_{uv}$. From these, we construct the assignments $A_{v}:\mathscr{G}_{1}^{E}\to\mathscr{G}_{1}$ and $B_{u}:\mathscr{G}_{1}^{D}\to\mathscr{G}_{1}$ by setting $A_{v}$ to be the $f_{V}(v)$-th projection for each $v\in V$, and $B_{u}$ to be the $f_{U}(u)$-th projection for each $u\in U$. Observe that the projections are folded over $\mathrm{id}_{\mathscr{H}_{1}}$, so $A_{v}=(A_{v})_{\mathrm{id}_{\mathscr{H}_{1}}}$. Moreover, note that, for a projection function $B:\mathscr{G}^{D}\to\mathscr{G}$, $\mathbf{b}\in\mathscr{G}^{D}$, and $s\in\{s_{1},s_{2}\}$, we have $B(\mathbf{b}^{s})^{s}=B(\mathbf{b})$. Then, for each edge $\{u,v\}$, each ${\mathbf{a}}\in\mathscr{G}_{1}^{E}$, ${\mathbf{b}},\bm{\nu}\in\mathscr{G}^{D}_{1}$ and ${\mathbf{c}}={\mathbf{b}}^{-1}({\mathbf{a}}\circ\pi_{uv})^{-1}\bm{\nu}$ we have
@@ 4.2 Soundness
$\left|\mathbb{E}_{uv,\mathbf{a}}\left[\mathrm{tr}\left(\left(\sum_{\tau\in\widehat{\mathscr{G}_{1}^{E}},\tau\neq 1}\sum_{s,t\in N_{\tau}}\dim_{\tau}\widehat{\mathcal{A}}(\tau_{s,t})\tau_{s,t}({\mathbf{a}})\right)\times\right.\right.\right.$
@@
$\left|\mathbb{E}_{uv,\mathbf{a}}\left[\mathrm{tr}\left(\left(\sum_{\tau\in\widehat{\mathscr{G}_{1}^{E}},\tau\neq 1}\sum_{s,t\in N_{\tau}}\dim_{\tau}\widehat{\mathcal{A}}(\tau_{s,t})\tau_{s,t}({\mathbf{a}})\right)\right.\right.\right.\times$
@@
$\left|\mathbb{E}_{uv,\mathbf{a}}\left[\mathrm{tr}\left(\left(\sum_{\tau\in\widehat{\mathscr{G}_{1}^{E}},\tau\neq 1}\sum_{s,t\in N_{\tau}}\dim_{\tau}\widehat{\mathcal{A}}(\tau_{s,t})\tau_{s,t}({\mathbf{a}})\right)\right.\right.\right.\times$
@@
$\left|\mathbb{E}_{uv,\mathbf{a}}\left[\mathrm{tr}\left(\left(\sum_{\tau\in\widehat{\mathscr{G}_{1}^{E}},\tau\neq 1}\sum_{s,t\in N_{\tau}}\dim_{\tau}\widehat{\mathcal{A}}(\tau_{s,t})\tau_{s,t}({\mathbf{a}})\right)\right.\right.\right.\times$
@@
Because $F$ is Hermitian, there must be a unitary matrix $U$ and a diagonal matrix $D$ satisfying $F=U^{*}DU$. We show that $F$’s eigenvalues are $1$, with multiplicity $\langle\chi_{\omega},\chi_{R_{\mathscr{G}_{2},\mathscr{H}_{2}}}\rangle$, and 00, with multiplicity $\dim_{\omega}-\langle\chi_{\omega},\chi_{R_{\mathscr{G}_{2},\mathscr{H}_{2}}}\rangle$. For each $\rho\in\widehat{\mathscr{H}_{2}}$, let $n_{\rho}$ be the multiplicity of $\rho$ in $\omega|_{\mathscr{H}_{2}}$. By Lemma 2.9, there is an invertible matrix $T$ such that $\omega(h)=T^{-1}(\bigoplus_{\rho\in\widehat{\mathscr{H}_{2}}}n_{\rho}\rho(h))T$ for all $h\in\mathscr{H}_{2}$. This way,
@@
[…]
@@
()^$\widehat{(\mathcal{B}*\mathcal{B})}_{y,x}(\rho_{i,j})=\sum_{z\in N_{\omega}}\sum_{k\in N_{\rho}}\widehat{{\mathcal{B}}_{y,z}}(\rho_{i,k})\widehat{{\mathcal{B}}_{z,x}}(\rho_{k,j}).$
@@
$\frac{\xi^{2}}{\dim_{\omega}^{6}}\leq$ $\mathbb{E}_{uv}\left[\sum_{\begin{subarray}{c}\rho\in\widehat{\mathscr{G}_{1}^{D}},|\rho|<\kappa,\\ i,j,k\in N_{\rho}\end{subarray}}\sum_{\begin{subarray}{c}\tau\in\widehat{\mathscr{G}_{1}^{E}},\tau\neq 1\\ \tau\sim\rho^{\pi},s,t\in N_{\tau}\end{subarray}}\dim_{\tau}\dim_{\rho}\left|(1-\epsilon)^{|\rho|}\langle\tau_{s,t},\rho^{\pi}_{j,i}\rangle\widehat{{\mathcal{B}}_{z,x}}(\rho_{k,j})\right|^{2}\right.$
×$\left.\times\sum_{\begin{subarray}{c}\rho\in\widehat{\mathscr{G}_{1}^{D}},|\rho|<\kappa,\\ i,j,k\in N_{\rho}\end{subarray}}\sum_{\begin{subarray}{c}\tau\in\widehat{\mathscr{G}_{1}^{E}},\tau\neq 1,\\ \tau\sim\rho^{\pi},s,t\in N_{\tau}\end{subarray}}\dim_{\tau}\dim_{\rho}\left|\widehat{\mathcal{A}_{x,y}}(\tau_{s,t})\widehat{{\mathcal{B}}_{y,z}}(\rho_{i,k})\right|^{2}\right].$ (14)
@@
[…]

### v3

@@ 1 Introduction
In this work, we strengthen the optimal inapproximability results for $\operatorname{3-LIN}$ over finite groups by establishing NP-hardness of beating the random assignment threshold even if the instance is almost satisfiable in an arbitrarily more restrictive setting. Formally, this is captured by fixing (not one but) two groups and a homomorphism between them, following the framework of promise CSPs [AGH17, BBKO21]. In detail, (decision) promise CSPs [BBKO21] can be seen as a qualitative form of approximation: Each constraint comes in two forms, a strong one and a weak one. The promise is that there is a solution satisfying all constraints in the strong form while the (potentially easier) goal is to find a solution satisfying all constraints in the weak form. An example of a strong vs. weak constraint on the same, say Boolean, domain is 11-in-33 vs NAE, where the former is $\{(0,0,1),(0,1,0),(1,0,0)\}$ and the latter is $\{(0,0,1),(0,1,0),(1,0,0),(1,1,0),(1,0,1),(0,1,1)\}$. NAE is weaker as the relation contains more tuples. While these two constraint relations capture the well-known NP-hard problems of 1-in-3-SAT and Not-All-Equal-SAT respectively [Sch78], finding an NAE-assignment turns out to be doable in polynomial time under the promise that a 1-in-3-assignment exists [BG21]! [^1] The algorithm involves solving an instance of $\operatorname{3-LIN}$ over the integers and rounding positive integers to $1$ and non-positive integers to 0, demonstrating the importance of $\operatorname{3-LIN}$ among promise CSPs. For constraints on different domains, the notion of strong vs. weak constraint is captured by a homomorphism between the (sets of all) constraint relations; in the example above, the homomorphism is just the identity function. The exact solvability of $\operatorname{3-LIN}$ in the promise setting was resolved in [LŽ25].
@@
Recent work of Barto et al. [BBK+24] initiated an algebraic investigation of (quantitative) approximation of promise CSPs. In the context of $\operatorname{3-LIN}$, here are two simple examples captured by this framework. First, let $\mathscr{G}$ be a group and $\mathscr{H}$ be a subgroup of $\mathscr{G}$. Given an almost-satisfiable system over the subgroup $\mathscr{H}$, maximise the number of satisfied equations over $\mathscr{G}$. Our results imply that beating the random assignment over $\mathscr{H}$ is NP-hard. In the second example, consider a group $\mathscr{G}$, a normal subgroup $\mathscr{H}$, and an almost-satisfiable system over $\mathscr{G}$. The goal this time is to maximise the number of satisfied equations in the system over the quotient $\mathscr{G}/\mathscr{H}$. Our results show that doing better than the random assignment over $\mathscr{G}/\mathscr{H}$ is NP-hard. More generally, going beyond subgroups and quotients of a given group, we fix two groups $\mathscr{G}_{1}$ and $\mathscr{G}_{2}$ and a group homomorphism $\varphi$ from a subgroup $\mathscr{H}_{1}$ of $\mathscr{G}_{1}$ to a subgroup $\mathscr{H}_{2}$ of $\mathscr{G}_{2}$ with the property that $\varphi$ extends to a group homomorphism from $\mathscr{G}_{1}$ to $\mathscr{G}_{2}$. Given an almost-satisfiable system of equations over $\mathscr{G}_{1}$ with constants in $\mathscr{H}_{1}$, the goal is to maximise the number of satisfied equations over $\mathscr{G}_{2}$ where the constants are interpreted in $\mathscr{H}_{2}$ via $\varphi$. Our main result establishes that doing better than the random assignment over $\mathscr{H}_{2}$ is NP-hard.
Theorem (Main theorem, informal statement).
Let $\mathscr{G}_{1}$ and $\mathscr{G}_{2}$ be finite groups, $\mathscr{H}_{1}$ a subgroup of $\mathscr{G}_{1}$, and $\varphi$ a group homomorphism from $\mathscr{H}_{1}$ to $\mathscr{H}_{2}=\textnormal{Im}(\varphi)$, which is a subgroup of $\mathscr{G}_{2}$. For arbitrary $\epsilon,\delta>0$, given a system of linear equations over $\mathscr{G}_{1}$ with constants in $\mathscr{H}_{1}$ that admits a solution satisfying at least a $(1-\epsilon)$-fraction of the constraints, it is NP-hard to find a solution over $\mathscr{G}_{2}$ with constants in $\mathscr{H}_{2}$ that satisfies at least a $(1/|\mathscr{H}_{2}|+\delta)$-fraction of the constraints, where the constants are interpreted via $\varphi$.
For a precise statement of the main result, cf. Theorem 3.1. Thus we give an optimal inapproximability result for a natural and fundamental fragment of promise CSPs, systems of linear equations.
@@
As our second main contribution, in Section 5 we relate our results to a recent theory of Barto et al. [BBK+24], who developed a systematic approach to study (in)approximability of promise CSPs, which includes approximability of promise linear equations, from the viewpoint of universal algebra. In particular, we show that the proof of our main result (Theorem 3.1) implies that the collection of symmetries222called the valued minion of plurimorphisms in [BBK+24]. of Promise $\operatorname{3-LIN}$ can be mapped homomorphically to the collection of symmetries of Gap Label Cover, a condition that, based on the algebraic theory from [BBK+24], is known to guarantee NP-hardness of the former problem. Thus, our result fits within the algebraic theory of Barto et al. [BBK+24].
@@
The structure of the paper is as follows. In Section 2 we set the notation and present the necessary technical background on Fourier analysis over non-Abelian groups. Section 3 presents the main results with discussion and also provides a sketch of the main proof: In Section 3.1 we present the reduction from Gap Label Cover to Promise $\operatorname{3-LIN}$, and in Section 3.2 we give an overview of the techniques used in the analysis of this reduction and of the main challenges that arise in extending previous work to the promise setting. Section 4 then gives all technical details of the proof of the main result, with the completeness analysis in Section 4.1 and the soundness analysis in Section 4.2. Finally, in Section 5 we show that our results fit within the theory of Barto et al. [BBK+24].
@@ 2 Background
A group consist of a finite set $\mathscr{G}$ equipped with a binary associative operation $\cdot\,$, a distinct element $1\in\mathscr{G}$ which acts as the identity of $\cdot\,$, and where each element $g\in\mathscr{G}$ has an inverse under $\cdot\,$, denoted by $g^{-1}$. When there is no ambiguity, we denote the group product $g\cdot h$ simply by $gh$. A group $\mathscr{G}$ is Abelian if $gh=hg$ for every $g,h\in\mathscr{G}$. A subset $\mathscr{H}\subseteq\mathscr{G}$ of a group $\mathscr{G}$ is called a subgroup of $\mathscr{G}$, denoted by $\mathscr{H}\leq\mathscr{G}$, if $\mathscr{H}$ equipped with the group operation of $\mathscr{G}$ forms a group. Given a group $\mathscr{G}$, a subgroup $\mathscr{H}$ of $\mathscr{G}$, and an element $g\in\mathscr{G}$, the right coset of $\mathscr{H}$ in $\mathscr{G}$ by $g$ is the set $\mathscr{H}g:=\{hg\mid h\in\mathscr{H}\}$. The set of right cosets of $\mathscr{H}$ in $\mathscr{G}$ is denoted $\mathscr{H}\backslash\mathscr{G}$. Let $N$ be a finite set. The $N^{\text{th}}$ direct power of $\mathscr{G}$, denoted by $\mathscr{G}^{N}$, is the group whose elements are $N$-tuples $\mathbf{g}\in\mathscr{G}^{N}$ of elements from $\mathscr{G}$, and where the group operation is taken component-wise, i.e., $\mathbf{g}\cdot\mathbf{h}(n)=\mathbf{g}(n)\cdot\mathbf{h}(n)$ for each $n\in N$. If $\mathscr{H}\leq\mathscr{G}$, we define $(h\cdot\mathbf{g})(n)=h\cdot\mathbf{g}(n)$ for each $h\in\mathscr{H}$ and $\mathbf{g}\in\mathscr{G}^{N}$. With this notation, the notion of coset extends to include the right cosets of $\mathscr{H}$ in $\mathscr{G}^{N}$ in a natural way.
@@ 2.1 Fourier Analysis over non-Abelian Groups
We define the right-regular representation of $\mathscr{G}$ on $\mathscr{H}\backslash\mathscr{G}$, denoted $R_{\mathscr{G},\mathscr{H}}$, to be the representation of $\mathscr{G}$ with $N_{R_{\mathscr{G},\mathscr{H}}}=\mathscr{H}\backslash\mathscr{G}$ whose entries are defined as follows. Given two cosets $\mathscr{H}g_{1},\mathscr{H}g_{2}$ and an element $g\in\mathscr{G}$, the $(\mathscr{H}g_{1},\mathscr{H}g_{2})$-entry of $R_{\mathscr{G},\mathscr{H}}(g)$ is $1$ if $\mathscr{H}g_{1}=\mathscr{H}g_{2}g^{-1}$, and zero otherwise. In other words, $R_{\mathscr{G},\mathscr{H}}(g)$ is the matrix describing the permutation on the cosets $\mathscr{H}\backslash\mathscr{G}$ resulting from right-multiplication with $g^{-1}$. The representation $R_{\mathscr{G},\mathscr{H}}$ is the representation induced on $\mathscr{G}$ by the trivial representation of $\mathscr{H}$. When $\mathscr{H}$ is the trivial subgroup containing just the identity element, we simply write $R_{\mathscr{G}}$ for $R_{\mathscr{G},\mathscr{H}}$. The representation $R_{\mathscr{G}}$ is known as the right-regular representation of $\mathscr{G}$. An application of the Frobenius Reciprocity Law yields the following result.
@@
Let $N$ be a finite set. Given a pair of functions function $F,H:\mathscr{G}\to\mathbb{C}^{N\times N}$, we define their convolution FHF*H by
@@ 2.1.1 Fourier Analysis over Direct Products
Let $\mathscr{G}$ be a finite group, $D$ a finite set, and $0<\epsilon<1$ a real number. Let $\bm{\nu}=(\nu_{d})_{d\in D}$ be a random element from $\mathscr{G}^{D}$ chosen as follows. Independently for each $d\in D$, the element $\nu_{d}$ equals $1_{\mathscr{G}}$ with probability $1-\epsilon$, and is picked uniformly at random from $\mathscr{G}$ otherwise. Let $F:\mathscr{G}^{D}\to\mathbb{C}$ and define $H(\mathbf{a}):=\mathbb{E}_{\bm{\nu}}[F(\mathbf{a}\cdot\bm{\nu})]$. Then, for each $\rho\in\widehat{\mathscr{G}^{D}}$ and $i,j\in N_{\rho}$, we have
@@ 3 Overview of Results
Let $(\mathscr{G}_{1},\mathscr{G}_{2},\varphi)$ be a cubic template and $0<s\leq c<1$. Then, $\operatorname{3-LIN}(\mathscr{G}_{1},\mathscr{G}_{2},\varphi,c,s)$ is tractable if $s\leq 1/|\textnormal{Im}(\varphi)|$ and NP-hard otherwise.
@@
Given a weighted system of equations over $(\mathscr{G}_{1},\mathscr{G}_{2},\varphi)$, consider its set of unsatisfiable equations. Since $\varphi$ extends to a full homomorphism, if the total weight of the set of unsatisfiable equations is more than 1c1-c, then the instance cannot be $c$-satisfiable in $\mathscr{G}_{1}$, hence, reject. Otherwise, the random assignment over $\textnormal{Im}(\varphi)$ satisfies at least a $1/|\textnormal{Im}(\varphi)|$-fraction of the satisfiable equations over $\mathscr{G}_{2}$, which is at least a $c/|\textnormal{Im}(\varphi)|$-fraction of the entire system. It is a simple corollary of Theorem 3.1 that this algorithm is optimal for non-cubic groups, leading to the following result. Details are deferred to Appendix A.
@@
Let $(\mathscr{G}_{1},\mathscr{G}_{2},\varphi)$ be a non-cubic template and $0<s\leq c<1$. Then, $\operatorname{3-LIN}(\mathscr{G}_{1},$ $\mathscr{G}_{2},\varphi,c,s)$ is tractable if $s/c\leq 1/|\textnormal{Im}(\varphi)|$ and NP-hard otherwise.
@@ 3.2 Proof Outline
[…]
@@ 4.1 Completeness
Let $f_{U}:U\to D$, $f_{V}:V\to E$ be assignments satisfying the original GLC instance $\Sigma$ - that is to say, these assignments are such that $\pi_{uv}(f_{U}(u))=f_{V}(v)$ for each edge $\pi_{uv}$. From these, we construct the assignments $A_{v}:\mathscr{G}_{1}^{E}\to\mathscr{G}_{1}$ and $B_{u}:\mathscr{G}_{1}^{D}\to\mathscr{G}_{1}$ by setting $A_{v}$ to be the $f_{V}(v)$-th projection for each $v\in V$, and $B_{u}$ to be the $f_{U}(u)$-th projection for each $u\in U$. Observe that the projections are folded over $\mathrm{id}_{\mathscr{H}_{1}}$, so $A_{v}=(A_{v})_{\mathrm{id}_{\mathscr{H}_{1}}}$. Moreover, note that, for a projection function $B:\mathscr{G}^{D}\to\mathscr{G}$, $\mathbf{b}\in\mathscr{G}^{D}$, and $s\in\{s_{1},s_{2}\}$, we have $B(\mathbf{b}^{s})^{s}=B(\mathbf{b})$. Then, for each edge $\{u,v\}$, each ${\mathbf{a}}\in\mathscr{G}_{1}^{E}$, ${\mathbf{b}},\bm{\nu}\in\mathscr{G}^{D}_{1}$ and ${\mathbf{c}}={\mathbf{b}}^{-1}({\mathbf{a}}\circ\pi_{uv})^{-1}\bm{\nu}$ we have
@@ 4.2 Soundness
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Because $F$ is Hermitian, there must be a unitary matrix $U$ and a diagonal matrix $D$ satisfying $F=U^{*}DU$. We show that $F$’s eigenvalues are $1$, with multiplicity $\langle\chi_{\omega},\chi_{R_{\mathscr{G}_{2},\mathscr{H}_{2}}}\rangle$, and 0, with multiplicity $\dim_{\omega}-\langle\chi_{\omega},\chi_{R_{\mathscr{G}_{2},\mathscr{H}_{2}}}\rangle$. For each $\rho\in\widehat{\mathscr{H}_{2}}$, let $n_{\rho}$ be the multiplicity of $\rho$ in $\omega|_{\mathscr{H}_{2}}$. By Lemma 2.9, there is an invertible matrix $T$ such that $\omega(h)=T^{-1}(\bigoplus_{\rho\in\widehat{\mathscr{H}_{2}}}n_{\rho}\rho(h))T$ for all $h\in\mathscr{H}_{2}$. This way,
@@
$\leq\$$\left|\mathrm{tr}\left(\sum_{\begin{subarray}{c}\rho\in\widehat{\mathscr{G}_{1}^{D}}\\ |\rho|\geq\kappa\end{subarray}}\dim_{\rho}(1-\epsilon)^{|\rho|}\mathcal{C}({\mathbf{a}},{\mathbf{a}},\rho))\right)\right|+\frac{1}{|\mathscr{G}_{1}^{E}|}\sum_{{\mathbf{g}}\in\mathscr{G}_{1}^{E}}\left|\mathrm{tr}\left(\sum_{\begin{subarray}{c}\rho\in\widehat{\mathscr{G}_{1}^{D}}\\ |\rho|\geq\kappa\end{subarray}}\dim_{\rho}(1-\epsilon)^{|\rho|}\mathcal{C}({\mathbf{g}},{\mathbf{a}},\rho))\right)\right|$
$\leq\$$2\mathrm{tr}\left(\sum_{\begin{subarray}{c}\rho\in\widehat{\mathscr{G}_{1}^{D}}\\ |\rho|\geq\kappa\end{subarray}}\sum_{i\in N_{\rho}}\dim_{\rho}(1-\epsilon)^{|\rho|}\widehat{(\mathcal{B}*\mathcal{B})}(\rho_{i,i})\right).$
@@
$\widehat{(\mathcal{B}*\mathcal{B})}_{y,x}(\rho_{i,j})=\sum_{z\in N_{\omega}}\sum_{k\in N_{\rho}}\widehat{{\mathcal{B}}_{y,z}}(\rho_{i,k})\widehat{{\mathcal{B}}_{z,x}}(\rho_{k,j}).$
@@
$\frac{\xi^{2}}{\dim_{\omega}^{6}}\leq\$$\mathbb{E}_{uv}\left[\sum_{\begin{subarray}{c}\rho\in\widehat{\mathscr{G}_{1}^{D}},|\rho|<\kappa,\\ i,j,k\in N_{\rho}\end{subarray}}\sum_{\begin{subarray}{c}\tau\in\widehat{\mathscr{G}_{1}^{E}},\tau\neq 1\\ \tau\sim\rho^{\pi},s,t\in N_{\tau}\end{subarray}}\dim_{\tau}\dim_{\rho}\left|(1-\epsilon)^{|\rho|}\langle\tau_{s,t},\rho^{\pi}_{j,i}\rangle\widehat{{\mathcal{B}}_{z,x}}(\rho_{k,j})\right|^{2}\right$.
$\left.\times\sum_{\begin{subarray}{c}\rho\in\widehat{\mathscr{G}_{1}^{D}},|\rho|<\kappa,\\ i,j,k\in N_{\rho}\end{subarray}}\sum_{\begin{subarray}{c}\tau\in\widehat{\mathscr{G}_{1}^{E}},\tau\neq 1,\\ \tau\sim\rho^{\pi},s,t\in N_{\tau}\end{subarray}}\dim_{\tau}\dim_{\rho}\left|\widehat{\mathcal{A}_{x,y}}(\tau_{s,t})\widehat{{\mathcal{B}}_{y,z}}(\rho_{i,k})\right|^{2}\right].$ (14)
@@
This inequality suggests the following randomised strategy to construct an assignment for the original instance $\Sigma$ of Gap Label Cover. Fix indices $x,y,z\in N_{\omega}$ as described above. Given a vertex $u\in U$, the assignment $d_{u}$ is chosen randomly by picking a representation $\rho\in\widehat{\mathscr{G}_{1}^{D}}$ with probability at least $\sum_{i,k\in N_{\rho}}\dim_{\rho}|\widehat{{\mathcal{B}}_{y,z}}(\rho_{i,k})|^{2}$, which is well-defined by Plancherel’s theorem and the fact that $\omega$ is unitary, and then picking $d_{u}$ uniformly at random among those elements $d\in D$ for which $\rho^{d}$ is non-trivial if such $d$ exists, and give up otherwise. Similarly, given a vertex $v\in V$, the assignment $e_{v}$ is chosen by picking a non-trivial representation $\tau\in\widehat{\mathscr{G}_{1}^{E}}$ with probability at least $\sum_{s,t\in N_{\tau}}\dim_{\tau}|\widehat{\mathcal{A}_{x,y}}(\tau_{s,t})|^{2}$, and then picking $e_{v}\in E$ uniformly at random among those elements $e\in E$ for which $\tau^{e}$ is non-trivial. This is also well-defined by Plancherel’s theorem and the fact that $\omega$ is unitary, and additionally, note that picking $e_{v}$ with $\tau^{e_{v}}$ non-trivial is always possible since (15) guarantees that there always exists some non-trivial $\tau\in\widehat{\mathscr{G}_{1}^{E}}$ with $|\widehat{\mathcal{A}_{x,y}}(\tau_{s,t})|>0$. The probability that this random assignment satisfies a given constraint $u\xrightarrow{\pi}v$ in the GLC instance is at least the joint probability of the following events: (1) the representations $\rho,\tau$ satisfy $|\rho|,|\tau|<\kappa$, and (2) the chosen element $e_{v}\in E$ satisfies $e_{v}=\pi(d_{u})$. Because of (15), this probability is at least $\xi^{2}\kappa^{-1}|\mathscr{G}_{1}|^{-\kappa}\dim_{\omega}^{-6}$. This strategy can be derandomised to obtain an assignment of $\Sigma$ satisfying at least the same fraction of its constraints, proving the result. ∎